# Prompt-Driven Building Footprint Extraction in Aerial Images with Offset-Building Model

### v3

@@ III-B3 Reference Offset Augment Module (ROAM)
In DETR[45], the box coder uses the size of input images to scale down the output results, enabling better fitting for the FFN. However, this strategy cannot be applied to the Offset Coder. This is because building offsets in the x and y directions often approach zero. If the Offset Coder copied encoding and decoding processes of DETR’s Box Coder entirely, the loss would become even closer to zero. To improve the offset quality, we propose ROAM. Each head in ROAM is similar except the scale factor of Offset Coder.
@@
Here, $\mathcal{O}_{R}^{i}$ represents the ith Reference Offset, $\mathcal{O}_{A}^{i}$ represents the ith Adaptive Offset, and $\omega^{i}$ is the indicator value, taking values of 0 or 1, indicating whether the corresponding Adaptive Offset affects the output $\mathcal{O}^{i}$. The average of the Reference Offset and $n$ Adaptive Offsets is the final output.
@@
where $\mathcal{L}_{ROAM}$ is the loss of ROAM, $\mathcal{L}_{roof}$ is CrossEntropy Loss of roof segmentation, and $\mathcal{L}_{building}$ is CrossEntropy Loss of building segmentation.
@@
$\mathcal{L}_{i}$ represents offset loss of the ith adaptive offset head. $\mathcal{L}_{B}$ represents offset loss from reference offset head. SmoothL1 Loss is applied for all offset heads.
@@ III-C Distance NMS and soft Distance NMS
Distance NMS (D-NMS) and soft Distance NMS (soft D-NMS) algorithms are specifically designed for refining building offset predictions. As shown in Table 1, big buildings often give out a more accurate offset in direction compared with shorter. Therefore, drawing on the concepts of NMS and soft NMS algorithms, we propose an NMS algorithm based on the predicted offset lengths.
D-NMS directly replaces directions of all offsets by the longest. But soft D-NMS is different, because it will using statistics to adjust the longer and shorter offsets.
@@ III-D ROI Prompt Based Offset Extraction
Figure 4: During the training of models, RPN (Region Proposal Network) is utilized to produce boxes. These boxes are employed with ROI Align to crop features, which is subsequently transformed into local features of identical size. These features are then inputted into multi-task heads for regression. In the inference phase, the RPN is deactivated, and models use manually provided boxes for ROI extraction.
@@ III-E Relative Height Map
With accurate roof segmentation and offset prediction, generating a relative height map is reliable. In the same image, all buildings will be allocate a relative height based on the length of their offsets. As shown in Fig. 5, the longest one will be defined as 1, at the inference stage, the top roof will be dragged to its footprint, and linearly interpolate the relative height.
@@
As the relationship between each building can be define by Eq. 6, the height map can be generated by finding a scale factor. This factor can be determined by directly measuring a height of a building in the given area, or be figured out by parameters of shooting camera and satellite.
@@ III-F1 Offset metrics
Referring COCO metric[61], we measure offsets based on different length group, and calculate the average of them. Above three losses, can described as,
@@ IV Experiment
Our OBM was trained on a server with 6 RTX 3090 for over 12 hours.
@@ IV-A Main result
TABLE 1: Main results of each box prompt models tested on BONAI[12].
Model 0,10 10,20 20,30 30,40 40,50 50,60 60,70 70,80 80,90 90,100 100$\infty$ $m$$\mathcal{L}$ $a$$\mathcal{L}$
$VL$ 5.48 4.34 6.39 6.80 8.41 11.8 15.6 20.3 17.2 18.2 53.9 15.4 6.12
LOFT $LL$ 4.15 2.75 4.51 5.18 6.52 8.70 12.2 16.7 12.9 14.4 49.9 12.6 4.51
$AL$ 0.55 0.21 0.17 0.11 0.10 0.14 0.15 0.16 0.13 0.13 0.15 0.18 0.32
Cas. LOFT $V$ 5.62 4.12 5.42 6.11 7.90 12.8 17.3 22.7 17.9 18.7 55.3 15.8 5.97
$LL$ 4.28 2.66 3.71 4.72 6.30 10.8 15.2 20.2 14.1 14.8 51.3 13.5 4.48
$AL$ 0.55 0.19 0.15 0.09 0.09 0.14 0.13 0.13 0.14 0.15 0.14 0.17 0.31
$VL$ 4.01 3.71 5.55 6.10 7.58 9.18 12.5 16.9 15.1 21.2 61.4 14.8 5.08
Ours $LL$ 3.20 2.55 4.14 4.96 6.14 7.77 10.8 15.9 13.8 19.5 60.1 13.5 4.03
$AL$ 0.37 0.15 0.14 0.08 0.08 0.07 0.08 0.06 0.06 0.10 0.11 0.12 0.22
Table 1 illustrates big performance difference between each model, but all models are inclined to predict better direction for buildings with longer offset. {$i,j$} in the first line represents the length range of ground offset. In the first 11 columns, we list out model performance in different length ranges. The performance of our model exceeds all other models. For buildings with offset over 90 pixels, OBM predicts a little worse than LOFT. This might be caused by the distribution of training data.
Table 2 shows DNMS algorithms still work well by correcting offset angles. While OBM already has a good ability to distinguish offset angles, the improvement of DNMS is limited. DNMS algorithms can also improve performance on the new dataset.
TABLE 2: Positive influence of DNMS algorithms on BOANI[12].
@@
In terms of mask ability, OBM still has a better performance compared with ROI prompt related models.
TABLE 3: IoU and Boundary IoU of predicted roofs on BONAI[12]
Model IoU Boundary IoU
LOFT 0.65 0.38
Cascade LOFT 0.68 0.40
Ours 0.73 0.43
TABLE 4: F1score of finally predicted footprints on BONAI[12]
Model LOFT Cas.LOFT Ours
F1score 0.495 0.504 0.581
As shown in Table 3, the roof regression of OBM is more accurate than LOFT and Cascade LOFT. In Fig. 7, some results were visualized. Footprint quality is highly relied on roof quality and offset accuracy. In Table 4, footprints of OBM are more accurate than other models.
@@ IV-B Results on generalization test
In this part, we newly annotated a dataset with over 7,000 samples. The spatial resolution of images is 0.5 m (same with BONAI[12]). All images are collected from Huizhou, China.
TABLE 5: Detail results on newly annotated dataset, without extra training
Model $mVL$ $mLL$ $mAL$ IoU BIoU
LOFT 10.40 9.009 0.770
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption LOFT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Cas. LOFT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Height(LOFT)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Height(Cas.LOFT)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Height(Ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Ground Truth
@@
Then we conduct inference on this dataset without extra training. The length of offsets includes plenty of information e.g. relative heights between each building. For better awareness of models, we visualized them in Fig. 7. Compared with former methods, our model has a 20.29% lower vector loss and a 76.36% lower angle loss in our brand-new dataset, shown in Table 5. Moreover, soft DNMS algorithm might not always provide better results than DNMS. This could attribute to the model performance: sometimes the direction of the longest offset is the best direction which has the minimal angle loss($AL$).
@@ V Ablation
LOFT and Cascade LOFT are originally designed for end-to-end instance segmentation. To clarify whether training in prompt like OBM would influence convergence, we design extra experiments in Table 7.
We find RPN can improve the offset quality of those models. Apart from above, using different tasks to lead the train may also influences final results. Model-related information, including FPS and parameters, is provided in Table 6.
TABLE 6: In this table, DNMS and soft DNMS correct the angle for those shorter offsets. The circumstances of OBM is similar.
Model FPS Params
Res50 + LOFT 3.20 77.70 M
Res50 + Cas. LOFT 2.50 180.74 M
SAM-b + OBM 1.69 90.40 M
In Table 8, We also provide different variants of OBM. Surprisingly, we find each variants has a considerable improvement compared with former methods.
TABLE 7: Ablation study of training methods. From this table, ROI based model needs to use RPN to reach better convergence, while our OBM needs segmentation tasks to lead training.

### v4

@@ III-B3 Reference Offset Augment Module (ROAM)
In DETR[52], the box coder uses the size of input images to scale down the output results, enabling better fitting for the FFN. However, this strategy cannot be applied to the Offset Coder. Based on our training experience, building offsets in the $x$ and $y$ directions often approach zero. If the Offset Coder copied encoding and decoding processes of DETR’s Box Coder entirely, the loss would become even closer to zero. We can deepen our understanding of this phenomenon through formulas in Eq. 2. The $\alpha$ is a constant in a box coder, and can be understood as a scale factor. According to statistics, most buildings have an offset length of less than 40 pixels. If $\alpha$ is too large, the most of $\mathcal{O}_{e}$ used in training will fluctuate around 0, and that means the model will learn nothing. Similarly, if $\alpha$ is too small, the value range of $\mathcal{O}_{e}$ will be extremely unstable, which will lead to unstable convergence of the model. However, larger $\alpha$ and smaller $\alpha$ are valuable for the learning of longer offsets and shorter offsets. To improve the offset quality, we proposed ROAM and combined aforementioned coders in adaptive heads. Each head in ROAM is similar except the scale factor of Offset Coder.
@@
Here, $\mathcal{O}_{R}^{i}$ represents the $i$ th Reference Offset, $\mathcal{O}_{A}^{i}$ represents the $i$ th Adaptive Offset, and $\omega^{i}$ is the indicator value, taking values of 0 or 1, indicating whether the corresponding Adaptive Offset affects the output $\mathcal{O}^{i}$. The average of the Reference Offset and $n$ Adaptive Offsets is the final output. The whole computing process can be described as: at the inference stage, Base Head, whose $\alpha$ in Offset Coder is moderate, will first give out a rough offset for each building. These offsets can be provided as references for Adaptive Heads, and, based on their lengths, they can be roughly divided into two groups: longer offsets and shorter offsets. Each group will use a different indicator to identify the offsets. Then, both groups will be assigned to different Adaptive Heads, and their assignments commonly follow the rule: longer offsets will use Adaptive Head with larger $\alpha$ and heads with smaller $\alpha$ for shorter offsets. With multi-heads outputs, the offsets will be finally determined. In Section V-B, we will introduce the detailed settings in our released model version.
@@
where $\mathcal{L}_{ROAM}$ is the loss of ROAM, $\mathcal{L}_{roof}$ is CrossEntropy Loss [71] of roof segmentation, and $\mathcal{L}_{building}$ is CrossEntropy Loss of building segmentation.
@@
$\mathcal{L}_{i}$ represents offset loss of the $i$ th adaptive offset head. $\mathcal{L}_{B}$ represents offset loss from reference offset head. SmoothL1 Loss [72] is applied for all offset heads.
@@ III-C Distance NMS and soft Distance NMS
Distance NMS (D-NMS) and soft Distance NMS (soft D-NMS) algorithms are specifically designed for refining building offset predictions. As shown in Table I, big buildings often give out a more accurate offset in direction compared with shorter. Therefore, drawing on the concepts of NMS and soft NMS algorithms, we propose an NMS algorithm based on the predicted offset lengths.
D-NMS directly replaces directions of all offsets by the longest. But soft D-NMS is different, because it will use statistics to adjust the longer and shorter offsets.
@@ III-D ROI Prompt Based Offset Extraction
Figure 4: During the training of models, RPN (Region Proposal Network) is utilized to produce boxes. These boxes are employed with ROI Align to crop features, which is subsequently transformed into local features of identical size. These features are then inputted into multitask heads for regression. In the inference phase, the RPN is deactivated, and models use manually provided boxes for ROI extraction.
@@ III-E Relative Height Map
With accurate roof segmentation and offset prediction, generating a relative height map is reliable. In the same image, all buildings will be allocated a relative height based on the length of their offsets. As shown in Fig. 5, the longest one will be defined as 1, at the inference stage, the top roof will be dragged to its footprint, and linearly interpolate the relative height.
@@
As the relationship between each building can be defined by Eq. 6, the height map can be generated by finding a scale factor. This factor can be determined by directly measuring a height of a building in the given area, or be figured out by parameters of shooting camera and satellite.
@@ III-F1 Offset metrics
Referring COCO metric[73], we measure offsets based on different length group, and calculate the average of them. Above three losses, can describe as,
@@ IV Experiment
TABLE I: Main results of each box prompt models tested on BONAI[15].
Model 0,10 10,20 20,30 30,40 40,50 50,60 60,70 70,80 80,90 90,100 100$\infty$ $m$$\mathcal{L}$ $a$$\mathcal{L}$ IoU BIoU
$VL$ 5.48 4.34 6.39 6.80 8.41 11.8 15.6 20.3 17.2 18.2 53.9 15.4 6.12
LOFT $LL$ 4.15 2.75 4.51 5.18 6.52 8.70 12.2 16.7 12.9 14.4 49.9 12.6 4.51 0.65 0.38
$AL$ 0.55 0.21 0.17 0.11 0.10 0.14 0.15 0.16 0.13 0.13 0.15 0.18 0.32
Cas. LOFT $VL$ 5.62 4.12 5.42 6.11 7.90 12.8 17.3 22.7 17.9 18.7 55.3 15.8 5.97
$LL$ 4.28 2.66 3.71 4.72 6.30 10.8 15.2 20.2 14.1 14.8 51.3 13.5 4.48 0.68 0.40
$AL$ 0.55 0.19 0.15 0.09 0.09 0.14 0.13 0.13 0.14 0.15 0.14 0.17 0.31
$VL$ 4.01 3.71 5.55 6.10 7.58 9.18 12.5 16.9 15.1 21.2 61.4 14.8 5.08
Ours $LL$ 3.20 2.55 4.14 4.96 6.14 7.77 10.8 15.9 13.8 19.5 60.1 13.5 4.03 0.73 0.43
$AL$ 0.37 0.15 0.14 0.08 0.08 0.07 0.08 0.06 0.06 0.10 0.11 0.12 0.22
@@
Our OBM was trained on a server with 6 RTX 3090 for over 12 hours. On each GPU, we have 1 sample; in other words, the Batch Size during training was 6. We choose SGD [76] as training optimizer, and the learning rate, momentum, and weight decay are 0.0025, 0.9 and 0.0001 respectively. GradClip [77] was used to ensure the model was trained on the right way at the beginning.
@@ IV-A Main result
The main results will focus on the models’ performance on BONAI dataset[15]. This dataset includes 3,000 and 300 images (shape 1024$\times$1024) for train-val and test respectively.
Table I illustrates big performance difference between each model, but all models are inclined to predict better direction for buildings with longer offset. {$i,j$} in the first line represents the length range of ground offset. In the first 11 columns, we list model performance in different length ranges. The performance of our model exceeds all other models. For buildings with offset over 90 pixels, OBM predicts a little worse than LOFT ($VL_{(90,100)}$ increased by 3 pixels). This might be caused by the distribution of training data. On comprehensive level, OBM dropped $mVL$, $mAL$, $aVL$ and $aA$ by 3.8%, 33.3%, 17.0% and 31.3% compared with prompt LOFT. Moreover, because OBM’s mask prediction ability was inherited from SAM, the roof performance of OBM significantly outperforms other models was within estimation. For example, roof IoU between predictions and Ground Truth (GT) improved by 12.3% compared with that of prompt LOFT.
Table II shows DNMS algorithms still work well by correcting offset angles. While OBM already has a good ability to distinguish offset angles, the improvement of DNMS is limited. DNMS algorithms can also improve performance on the new dataset. DNMS and soft DNMS model can correct $aAL$ by 0.04-0.44 for different models. The improvements are related with the model’s direction performance. e.g. OBM has a very good directional perception, improvements made by correcting angles using DNMS and soft DNMS are limited ($mAL$ and $aAL$ dropped only around 0.01). While these improvements on prompt LOFT and prompt Cascade LOFT are very significant, e.g. soft DNMS dropped $mAL$ and $aAL$ by 22.2% and 28.1% for prompt LOFT, and consequently $mVL$ and $aVL$ improved.
The improvements seem different from the huge gap between NMS and soft NMS, a supervising result can be found comparing Table II and Table IV: soft DNMS algorithm cannot always defeat the performance of DNMS. To understand this, we need to know the difference of their mechanisms. Soft NMS allows models to output more bounding boxes with low confidence, and end-to-end models are inclined to predict only a few number of valuable bounding boxes with higher scores. In other words, soft NMS can help the models recall more correct bounding boxes. However, soft DNMS and DNMS are different. Both methods mainly operate the direction of predicted offsets to correct results. Obviously, this process will not delete any bounding boxes. The final results will only be determined by the global performance and the performance of the tallest building respectively. Thus, there is a possibility that the longest offsets outperforms the global results in terms of a certain dataset at the most of the time. That is the reason why soft DNMS cannot always defeat DNMS.
TABLE II: Positive influence of DNMS algorithms on BOANI[15].
@@
In terms of mask ability, OBM still has a better performance compared with ROI prompt related models. In Table III, our model is compared with prompt-level LOFT, prompt-level Cascade LOFT, MTBR-Net(ICCV 2021)[50], LOFT(TPAMI 2023)[15] and MLS-BRN(CVPR 2024)[17] in terms of the ultimate footprint masks. The first half of the form demonstrates our model outperforms all prompt-level models and the second half denotes our model is better than end-to-end models on BONAI dataset. Comparing models in different mode, easily we discover prompt-level models can give more accurate footprint segmentations, but we need to emphasize that this comparison has limitations. Even for the same model LOFT, in prompt mode, LOFT will be told all potential buildings’ location, and the f1score of prompt LOFT improved by 15.1% compared with end-to-end LOFT. The results are comprehensible: end-to-end outputs need to consider the problem of mistake detection and omitted buildings, while prompt-level models utilize inputted boxes ensure all buildings can get a relatively correct answer.
TABLE III: F1score, Recall and Precision of finally predicted footprints on BONAI among different models[15]
Model F1score Recall Precision
prompt LOFT 0.740 0.853 0.676
prompt Cas.LOFT 0.775 0.867 0.718
Ours(OBM) 0.813 0.832 0.808
LOFT[15] 0.643 0.653 0.634
MTBR-Net[50] 0.636 0.629 0.643
MLS-BRN[17] 0.664 0.668 0.659
As shown in Table I, the roof regression of OBM is more accurate than LOFT and Cascade LOFT. In Fig. 8, some results were visualized. Footprint quality is highly relied on roof quality and offset accuracy.
@@ IV-B Results on generalization test
In this part, we newly annotated a dataset with over 7,000 instances. The spatial resolution of images (shape 1024$\times$1024) is 0.5 m (same with BONAI[15]). All images are collected from Huizhou, China (Coordinates: 23.1125° N, 114.4155° E). Huizhou is a coastal city of Guangdong Province. Due to its advantageous geographical location, the reform and opening-up policy has rapidly transformed Huizhou into a modern city. This allowing Huizhou to have many complex buildings, including bungalows and skyscrapers. Within huizhou test set, 43.1% of building offsets are longer than 10 pixels. Fig. 7 displayed some figures in this test dataset.
Figure 7: Some samples of Huizhou test dataset. In given pictures, red and green boundaries represent the edges of roof masks and footprint masks.
TABLE IV: Detail results on newly annotated dataset, without extra training
Model $aVL$ $aLL$ $aAL$ IoU BIoU
p. LOFT 10.40 9.009 0.770
@@
Then we conduct inference on this dataset without extra training. The length of offsets includes plenty of information e.g. relative heights between each building. For better awareness of models, we visualized them in Fig. 8. Compared with former methods, our model has a 20.29% lower vector loss and a 76.36% lower angle loss in our brand-new dataset, shown in Table IV. Moreover, soft DNMS algorithm might not always provide better results than DNMS. This could attribute to the model performance: sometimes the direction of the longest offset is the best direction which has the minimal angle loss($AL$).
LOFT
Cas. LOFT
Ours
Height(LOFT)
Height(Cas.LOFT)
Height(Ours)
Ground Truth
@@
To ensure the ability of our model, experiments were conducted on OmniCity-view3[78], and related results can be found on Table V. OmniCity-view3 is a dataset with offset annotations for 17,092 and 4,929 images (shape 512$\times$512) from train-val and test set respectively. BONAI has 165,881 annotations for training and 66.8% of them are longer than 10 pixels. In OmniCity-view3, there are 191,470 annotations, but 47.6% of them are shorter than 10 pixels.
In this experiment, we found a similar result with experiment on BONAI. Prompt-level test results widely outperform those end-to-end models. The reason may also be similar. Note that all the experimental models are trained on full OmniCity-view3 dataset. Our OBM enhances the F1-score by 5.7% relative to prompt LOFT. The results indicate that our approach betters the $aAL$ by 0.884 in comparison to prompt LOFT.
TABLE V: Results about offset and footprint quality on OmniCity-view3[78] dataset. OBM* represents model trained on BONAI’s zero-shot performance.
Model $aVL$ $aLL$ $aAL$ Precision Recall F1score
prompt LOFT 7.575 5.287 0.699 0.729 0.846 0.770
prompt Cas.LOFT 7.248 5.116 0.690 0.756 0.864 0.797
OBM(Ours) 6.691 5.145 0.636 0.785 0.861 0.814
OBM* 9.622 7.214 1.274 0.781 0.794 0.781
LOFT[15] - - - 0.688 0.722 0.705
MLS-BRN[17] - - - 0.696 0.751 0.723
Table VI illustrates DNMS algorithms can improve the global performance on OmniCity-view3[78]. The improvements on LOFT and Cascade LOFT are very clear, e.g. soft DNMS can reduce $aVL$ by 0.348 pixels (4.6%) and 0.369 pixels (5.1%) respectively. This table also implies that OBM has a very stable ability to give out the right direction of building offsets, so the improvement caused by DNMS algorithms is limited but exists.
TABLE VI: Results about offset correction with DNMS algorithms on OmniCity-view3 dataset.
Model $mVL$ $mLL$ $mAL$ $aVL$ $aLL$ $aAL$
prompt LOFT 54.31 48.53 0.652 7.575 5.287 0.699
+DNMS 54.59 48.53 0.663 7.563 5.287 0.668
+soft DNMS 54.14 48.66 0.623 7.227 5.301 0.604
prompt Cas.LOFT 52.90 48.37 0.616 7.248 5.116 0.690
+DNMS 53.15 48.37 0.641 7.216 5.116 0.641
+soft DNMS 52.84 48.48 0.596 6.879 5.123 0.583
OBM(Ours) 56.90 53.67 0.659 6.691 5.145 0.636
+DNMS 56.92 53.67 0.677 6.691 5.145 0.623
+soft DNMS 56.89 53.68 0.662 6.681 5.156 0.624
On this dataset, $mVL$ and $mLL$ are abnormal compared with that of BOANI. A contributing factor for this was the imbalance of data distribution, which finally makes the mistake prediction of those longer offsets (over 100 pixels). $m\mathcal{L}$ reflects more on the comprehensive performance of the model at various lengths of the predicted offset, and $a\mathcal{L}$ reflects the performance among the whole dataset. In the training set of OmniCity, there are 365,168 instance annotations among 17,092 images, but only 1,735 of them (0.475%) longer than 100 pixels. This makes $VL_{(100,\infty)}$ reach 322.9 pixels, and finally leads to the abnormality of $mVL$.
@@ V Ablation
V-A Model performance influenced by training strategy
LOFT and Cascade LOFT are originally designed for end-to-end instance segmentation. To clarify whether training in prompt like OBM would influence convergence, we design extra experiments in Table VII. From this table, we discover whether using RPN in training are a key component for final performance. e.g. LOFT with RPN can lower $mVL$ for over 41.06% compared with the LOFT trained without RPN. Apart from above, using different tasks to lead the train may also influence final results.
TABLE VII: Ablation study of training methods. From this table, ROI based model needs to use RPN to reach better convergence, while our OBM needs segmentation tasks to lead training.
@@ V-A Model performance influenced by training strategy
Model-related information, including FPS and parameters, is provided in Table VIII. FPS was computed on the same server with one RTX 3090. The size of OBM is small enough; however, transformer blocks featured with their lower efficiency. However, CNN based model can tolerate larger batch size while inferencing. Based on our experiments, the FPS of OBM was influences by the number of box prompts, while CNN based models perform more stably. For example, on BONAI dataset, the scale of images are larger than that of OmniCity, and the number of prompts are relatively larger. The speed of OBM was the slowest as shown in Table VIII. But on OmniCity, OBM is the fastest model.
TABLE VIII: This table illustrates the number of parameters in different model, and their computation effeciencies.
Model FPS Params
Res50 + LOFT 3.20 77.70 M
Res50 + Cas. LOFT 2.50 180.74 M
SAM-b + OBM 1.69 90.40 M
V-B Structure test of OBM
In Table X, We also provide different variants of OBM. Surprisingly, we find all variants has a considerable improvement compared with former methods.
To further the understanding of ROAM structure, we provide testing results of OBM with different number of Adaptive Head. The baseline OBM is one Base Head (BoxCoder $\alpha=200$) with three Adaptive Head (BoxCoder $\alpha=150,300,400$). In Table IX, we tested beseline OBM’s Base Head with different combination of Adaptive Heads. Although a simple combination always gets a worse answer, three-heads-assisting system finally gave out the best answer. The model was also retrained with only Base Head, shown in (-, -, -), which may imply ROAM system can inhance the model’s ability on offset predicting while training. To ensure which was the contributing factor to the improvement, we have calculated the accuracy of length prediction for different heads. The accuracy illustrates smaller BoxCoder $\alpha$ can better predict shorter offset, and the larger $\alpha$ can provide length compensation in the prediction. Worth to mention, all heads can give out an almost-same-good offset direction.
TABLE IX: Ablation study of ROAM on BONAI[15]. In the column of Model, we use (1,0,0) to represent base head with the first adaptive head. At last, we retrain an OBM with no adaptive head, and use (-,-,-) to denote this experiment.
Model $mVL$ $mLL$ $mAL$ $aVL$ $aLL$ $aAL$
(0,0,0) 14.93 13.61 0.117 5.118 4.051 0.221
(1,0,0) 14.91 13.56 0.117 5.149 4.076 0.220
(0,1,0) 24.76 23.55 0.118 10.27 9.416 0.221
(0,0,1) 39.09 38.11 0.118 15.01 14.26 0.222
(1,1,0) 14.65 13.34 0.116 5.115 4.054 0.219
(1,0,1) 14.23 13.48 0.116 5.978 4.941 0.219
(0,1,1) 22.26 21.02 0.117 9.283 8.408 0.221
(1,1,1) 14.84 13.53 0.117 5.083 4.027 0.220
(-, -, -) 17.40 15.91 0.138 5.651 4.422 0.256
V-C The robustness of the model to prompt errors